If 100 ml of a 0.5 M HCl solution are poured into 10 liters of water , what is the pOH
1
Expert's answer
2017-10-16T15:05:06-0400
pH+pOH=14
pH=-log([H^+ ])
HCl→H^++〖Cl〗^-
ν(HCl)=c (HCl)∙V(HCl)=0.1∙0.5=0.05 mol.
c〖(HCl)〗_(new solution)=ν(HCl)/V(HCl) =0.05/10.1=0.00495mol/l=c (H^+ ).
pH=2.31 ⇒pOH=11.69~11.7.
Answer
pOH=11.7.